Exercise 3-4

The following trial balance of Teal Co. does not balance.

TEAL CO.
TRIAL BALANCE
JUNE 30, 2017



Debit


Credit

Cash



$3,099
Accounts Receivable
$3,460



Supplies
1,029



Equipment
4,029



Accounts Payable



2,895
Unearned Service Revenue
1,429



Common Stock



6,229
Retained Earnings



3,229
Service Revenue



2,609
Salaries and Wages Expense
3,629



Office Expense
1,169



Totals

$14,745



$18,061



Each of the listed accounts should have a normal balance per the general ledger. An examination of the ledger and journal reveals the following errors.

1.
Cash received from a customer on account was debited for $570, and Accounts Receivable was credited for the same amount. The actual collection was for $750.
2.
The purchase of a computer printer on account for $729 was recorded as a debit to Supplies for $729 and a credit to Accounts Payable for $729.
3.
Services were performed on account for a client for $890. Accounts Receivable was debited for $890 and Service Revenue was credited for $89.
4.
A payment of $294 for telephone charges was recorded as a debit to Office Expense for $294 and a debit to Cash for $294.
5.
When the Unearned Service Revenue account was reviewed, it was found that service revenue amounting to $554 was performed prior to June 30 (related to Unearned Service Revenue).
6.
A debit posting to Salaries and Wages Expense of $899 was omitted.
7.
A payment on account for $206 was credited to Cash for $206 and credited to Accounts Payable for $260.
8.
A dividend of $804 was debited to Salaries and Wages Expense for $804 and credited to Cash for $804.


Prepare a correct trial balance.TEAL CO. TRIAL BALANCE JUNE 30, 2017 Credit Debit Cash $ $ Accounts Receivable Supplies Equipment Accounts Payable Unearned SAccounts Payable Unearned Service Revenue Common Stock Retained Earnings Service Revenue Salaries and Wages Expense Office Ex
Answer #1
Debit Credit
Cash 2691 =3099+(750-570)-294-294
Accounts Receivable 3280 =3460-(750-570)
Supplies 300 =1029-729
Equipment 4758 =4029+729
Accounts Payable 2429 =2895-260-206
Unearned Service Revenue 875 =1429-554
Common Stock 6229
Retained Earnings 3229
Dividends 804
Service revenue 3964 =2609+(890-89)+554
Salaries and wages expense 3724 =3629+899-804
Office expenses 1169
Totals 16726 16726